# Modeling and computation of the effective elastic behavior of parallelogram origami metamaterials

### v2

@@ 1 Introduction
Figure 1: Examples of soft modes of deformation in paper models of Miura and Eggbox origami. (a-b) Pinching and bending Miura Origami. (c-d) Analogous examples for Eggbox origami.
@@
This paper is organized as follows. Section 2 describes the geometry of parallelogram origami and summarizes the effective theory derived in [64]. Section 3 introduces our generalized elastic continuum model, and its equilibrium equations and boundary conditions. Section 4 provides the finite element formulation and discusses its implementation. Section 5 deals with examples and Section 6 concludes the work.
@@ 2.2 Coarse-graining rules and an effective description of origami soft modes
Figure 3: Illustration of the locally mechanistic coarse-graining rule. (a) A finely patterned parallelogram origami, with cells of size $\sim\ell$, covers a unit square domain. (b) A soft mode of this pattern corresponds to a spatially varying local mechanism.
Beyond the pure mechanisms, parallelogram origami can bend, twist and actuate in complex nonuniform ways at very little elastic energy by localizing most of the deformation at the folds. We call such deformations soft modes. As anyone who has folded a Miura origami knows, soft modes are the generic response of these patterns to loads. Here, we describe a theory that captures the effective behavior of all possible soft modes in parallelogram origami. As explained below, it is built on two coarse-graining rules that link the design and kinematics of the origami cells explicitly to its effective cell-averaged deformation.
The first coarse-graining rule originates from our earlier work on planar kirigami [67, 68]. Soft modes are locally mechanistic. At the scale of each unit cell they look like a mechanism. However, their features vary slowly from cell to cell. This rule reveals a fundamental relationship between the first fundamental form of the effective deformation $\mathbf{y}(\mathbf{x})$ of a soft mode
@@
and the cell-by-cell actuation $\theta=\theta(\mathbf{x})$, which we now think of as a spatially varying field. Fig. 3 illustrates these ideas by way of a concrete example. A reference parallelogram origami pattern is overlaid onto a square domain $\Omega=(0,1)^{2}$ with unit cells now of characteristic length $\sim\ell$ in Fig. 3(a) and subject to a soft mode of deformation in Fig. 3(b). Notice that, while the origami surface oscillates due to the underlying microstructure of the unit cell, corresponding vertices in the reference pattern (the ones related by a translation $m\ell\mathbf{u}_{0}+n\ell\mathbf{v}_{0}$, $m,n\in\mathbb{Z}$) evolve smoothly from cell to cell. Since the soft mode also exhibits the same corrugated features as the mechanism, this evolution is quantified by an underlying effective deformation $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$, rotation field $\mathbf{R}\colon\Omega\rightarrow SO(3)$, and actuation field $\theta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$ that satisfy (see Fig. 3)
$\mathbf{y}(\mathbf{x}+\ell\tilde{\mathbf{u}}_{0})-\mathbf{y}(\mathbf{x})\approx\ell\partial_{\mathbf{u}_{0}}\mathbf{y}(\mathbf{x})\approx\ell\mathbf{R}(\mathbf{x})\mathbf{u}(\theta(\mathbf{x})),$ (10)
$\mathbf{y}(\mathbf{x}+\ell\tilde{\mathbf{v}}_{0})-\mathbf{y}(\mathbf{x})\approx\ell\partial_{\mathbf{v}_{0}}\mathbf{y}(\mathbf{x})\approx\ell\mathbf{R}(\mathbf{x})\mathbf{v}(\theta(\mathbf{x})).$
It immediately follows that the fields $\theta$ and $\mathbf{I}(\mathbf{y})$ are coupled through the shape tensor in (8):
$\mathbf{I}(\mathbf{y})=\mathbf{A}^{T}(\theta)\mathbf{A}(\theta).$ (11)
While most of the deformation is localized at the folds in a soft mode, the panels must bend slightly to accommodate a non-uniform actuation from cell to cell. Our second coarse-graining rule is to constrain the effective theory by how neighboring unit cells with slightly bent panels fit together. This rule reveals a fundamental relationship between the actuation $\theta(\mathbf{x})$ and the second fundamental form of the effective deformation
@@
Fig. 4 illustrates the idea. Zoom into the local neighborhood of cells in a soft mode and identify the rotation of the base cell by $\mathbf{R}(\mathbf{x})$, as indicated, to be consistent with (LABEL:eq:keyCoarseGrain1) and Fig. 3. Given the slightly bent nature of all the panels in this description, the cells rotate relative to each other by a small amount. We quantify this kinematics at leading order by noticing that the rotation of the cell to the right and above the base cell satisfy
$\mathbf{R}(\mathbf{x}+\ell\tilde{\mathbf{u}}_{0})\approx\mathbf{R}(\mathbf{x})\big{(}\mathbf{I}+\ell(\bm{\omega}_{\mathbf{u}_{0}}(\mathbf{x})\times)\big{)},$ (13)
$\mathbf{R}(\mathbf{x}+\ell\tilde{\mathbf{v}}_{0})\approx\mathbf{R}(\mathbf{x})\big{(}\mathbf{I}+\ell(\bm{\omega}_{\mathbf{v}_{0}}(\mathbf{x})\times)\big{)},$
@@
As a final bit of manipulation, this constraint can be rewritten using the implied identity $\partial_{\mathbf{u}_{0}}\big{[}\mathbf{R}\mathbf{v}(\theta)\big{]}=\partial_{\mathbf{v}_{0}}\big{[}\mathbf{R}\mathbf{u}(\theta)\big{]}$ from (LABEL:eq:keyCoarseGrain1). Doing so eliminates the bending measures $\bm{\omega}_{\mathbf{u}_{0}}$, $\bm{\omega}_{\mathbf{v}_{0}}$ in favor of the second fundamental form $\mathbf{II}(\mathbf{y})$, giving an equivalent constraint in terms of the fields $\mathbf{y}$ and $\theta$:
@@ 2.3 Asymptotic analysis and an effective plate theory
The bulk of our effort in [64] goes into justifying the constraints in (11) and (15). We showed that every sufficiently smooth solution to these effective compatibility equations is the limit of a sequence of soft modes — origami deformations with negligible panel strain that approximate the effective deformation. The precise result is as follows: Let $\Omega\subset\mathbb{R}^{2}$ be a simply connected domain of characteristic length $\sim 1$, and let $\{\mathbf{x}_{i,j}^{\ell}\}\subset\mathbb{R}^{3}$ label the vertices of a reference parallelogram origami domain overlaid onto $\Omega$ whose cells are of characteristic length $\ell\ll 1$ (see, e.g., Fig. 3). Then, for any sufficiently smooth $\theta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$ and $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$ that solve (11) and (15), there is a corresponding deformation of the origami vertices $\{\mathbf{x}_{i,j}^{\ell}\}\rightarrow\{\mathbf{y}_{i,j}^{\ell}\}$ such that
$\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i+1,j}^{\ell}|-|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i,j+1}^{\ell}|-|\mathbf{x}_{i,j+1}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}$ (16)
$\qquad\qquad+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i+1,j}^{\ell}|-|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i,j+1}^{\ell}|-|\mathbf{x}_{i,j+1}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}=O(\ell^{2}),$
$|\mathbf{y}(\tilde{\mathbf{x}}_{i,j}^{\ell})-\mathbf{y}_{i,j}^{\ell}|=O(\ell)$
for all $(i,j)$, where $\tilde{\mathbf{x}}_{i,j}^{\ell}$ is the projection of $\mathbf{x}_{i,j}^{\ell}$ onto $\Omega$. The first statement in (LABEL:eq:mainResult) says that the strain of every panel is at most $\sim\ell^{2}$. The second says that the origami deformation approximates the effective deformation.
In fact, we can say more. Consider a protypical bar-and-hinge energy [26, 33] composed of stretching, bending and folding terms, denoted $E_{\text{tot}}^{\ell}(\{\mathbf{y}_{i,j}^{\ell}\})=E^{\ell}_{\text{str}}(\{\mathbf{y}_{i,j}^{\ell}\})+E^{\ell}_{\text{bend}}(\{\mathbf{y}_{i,j}^{\ell}\})+E^{\ell}_{\text{fold}}(\{\mathbf{y}_{i,j}^{\ell}\})$, and scale the moduli of these three terms by powers of $\ell$ such that “stretching stiffness” $\gg$ “bending stiffness” $\gg$ “folding stiffness”. With the help of the constructions above, we obtain the limit of this energy
@@
where $\mathbb{K}(\theta)\in\mathbb{R}^{3\times 2\times 3\times 2}$ is an explicit fourth order tensor with major symmetry $[\mathbb{K}(\theta)]_{i\alpha j\beta}=[\mathbb{K}(\theta)]_{j\beta i\alpha}$ that depends kinematically only on the actuation field $\theta$ (see [64] for the full definition). The set $A_{\text{eff}}$ denotes the class of admissible effective deformations and actuation fields describing a soft mode

### v3

@@ 1 Introduction
Figure 1: Examples of soft modes of deformation in paper models of Miura and Eggbox origami. (a-b) Pinching and bending in Miura Origami. (c-d) Analogous examples for Eggbox origami.
@@
This paper is organized as follows. Section 2 describes the geometry of parallelogram origami and summarizes the effective theory derived in [64]. Section 3 introduces our generalized elastic continuum model, and its equilibrium equations and boundary conditions. Section 4 provides the finite element formulation and discusses its implementation. Section 5 deals with examples and Section 6 concludes the work. Throughout, we make liberal use of tensor notation for conciseness. Vectors are indicated by boldface lowercase letters ($\mathbf{a}\in\mathbb{R}^{m}$), second-order tensors by boldface uppercase letters ($\mathbf{A}\in\mathbb{R}^{m\times n}$), third-order tensors by calligraphic uppercase letters ($\mathcal{A}\in\mathbb{R}^{m\times n\times p}$), and fourth-order tensors by “blackboard bold” uppercase letters ($\mathbb{A}\in\mathbb{R}^{m\times n\times p\times q}$).
@@ 2.2 Coarse-graining rules and an effective description of origami soft modes
Figure 3: Illustration of the locally mechanistic coarse-graining rule. (a) A finely patterned parallelogram origami, labeled $\{\mathbf{x}_{i,j}^{\ell}\}$, has cells of size $\sim\ell$ on an overall domain of size $\sim 1$. (b) A soft mode of this pattern deforms the vertices $\{\mathbf{x}_{i,j}^{\ell}\}\mapsto\{\mathbf{y}_{i,j}^{\ell}\}$ consistent with a spatially varying local mechanism. (c) The effective midsurface of the origami, denoted $\Omega$, deforms smoothly to $\mathbf{y}(\Omega)$ under a soft mode. (d) The tangent vectors induced by the effective deformation match the locally mechanistic response in (b).
Beyond the pure mechanisms, parallelogram origami can bend, twist and actuate in complex nonuniform ways at very little elastic energy by localizing most of the deformation at the folds. We call such deformations soft modes. As anyone who has folded a Miura origami knows, soft modes are the generic response of these patterns to loads. Here, we describe a theory that captures the effective behavior of all possible soft modes in parallelogram origami.
Let’s begin with some notation. A typical parallelogram origami metamaterial is a pattern of $M\times N$ repeating unit cells of length $|\mathbf{u}_{0}|$ and width $|\mathbf{v}_{0}|$ as in Fig. 2(c). Our starting point for coarse graining these metamaterials is shown in Fig. 3(a-b). We non-dimensionalize by fixing the width of the reference configuration of the origami pattern as $1$, thus introducing a characteristic cell size $\ell=1/N$. The vertices of this reference configuration are labeled by $\{\mathbf{x}_{i,j}^{\ell}\}\subset\mathbb{R}^{3}$, which indexes them by $i$ along the length and $j$ along the width. A corresponding soft mode of deformation takes each point $\mathbf{x}_{i,j}^{\ell}$ to a point $\mathbf{y}_{i,j}^{\ell}$ in 3D space, and is denoted by $\{\mathbf{y}_{i,j}^{\ell}\}\subset\mathbb{R}^{3}$. In [64], we showed that soft modes admit a continuum description quantified by the cell size $\ell$. Fig. 3(c-d) illustrates the main idea. At a basic level, we seek to replace the pattern $\{\mathbf{x}_{i,j}^{\ell}\}$ by a planer continuum domain $\Omega\subset\mathbb{R}^{2}$, reflecting its “midsurface”, with the goal of approximating the gross shape change of its soft modes $\{\mathbf{y}_{i,j}^{\ell}\}$ by continuum deformations of this surface $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$. We call such deformations effective deformations from hereon. As explained below, this continuum approach is made meaningful through two coarse-graining rules that link the design and kinematics of the origami’s cells explicitly to its effective deformations.
The first coarse-graining rule originates from our earlier work on planar kirigami [67, 68]. Soft modes are locally mechanistic. At the scale of each unit cell they look like a mechanism. However, their features vary slowly from cell to cell. This rule reveals a fundamental relationship between the first fundamental form of the effective deformation of a soft mode
@@
and the cell-by-cell actuation $\theta=\theta(\mathbf{x})$, which we now think of as a spatially varying field depending on $\mathbf{x}=(x_{1},x_{2})\in\Omega$. Here and throughout, the gradient operator $\nabla$ denotes differentiation with respect to the planar domain $\Omega$; thus $\nabla\mathbf{y}\in\mathbb{R}^{3\times 2}$ and $\mathbf{I}(\mathbf{y})\in\mathbb{R}^{2\times 2}_{\text{sym}}$.
To explain the coarse-graining rule, the vertex labeled $\mathbf{x}_{i,j}^{\ell}$ in Fig. 3(a) corresponds to a material point $\mathbf{x}\in\Omega$ in Fig. 3(c). A soft mode deforms this origami vertex to $\mathbf{y}_{i,j}^{\ell}$, as shown in Fig. 3(b), which corresponds to a spatial point $\mathbf{y}(\mathbf{x})$ on the effective surface in Fig. 3(d). Notice that, while the origami surface oscillates due to the underlying microstructure of the unit cell, the deformation of every other vertex $\mathbf{x}_{i+2m,j+2n}^{\ell}\mapsto\mathbf{y}_{i+2m,j+2n}^{\ell}$ is consistent with a smooth underlying surface. As the corrugated features are also consistent locally with that of a perfect mechanism, we can quantify the tangent planes of this surface by a rotation field $\mathbf{R}\colon\Omega\rightarrow SO(3)$ and actuation field $\theta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$ that satisfy
$\mathbf{y}_{i+2,j}^{\ell}-\mathbf{y}_{i,j}^{\ell}\approx\ell\mathbf{R}(\mathbf{x})\mathbf{u}(\theta(\mathbf{x})),\qquad\mathbf{y}_{i,j+2}^{\ell}-\mathbf{y}_{i,j}^{\ell}\approx\ell\mathbf{R}(\mathbf{x})\mathbf{v}(\theta(\mathbf{x}))$ (10)
(compare Fig. 2(c-d) to Fig. 3(a-b)). On the other hand, as a matter of standard continuum mechanics, we have that $\mathbf{y}$ in Fig. 3(c-d) satisfies
$\mathbf{y}(\mathbf{x}+\ell\tilde{\mathbf{u}}_{0})-\mathbf{y}(\mathbf{x})\approx\ell\partial_{\mathbf{u}_{0}}\mathbf{y}(\mathbf{x}),\qquad\mathbf{y}(\mathbf{x}+\ell\tilde{\mathbf{v}}_{0})-\mathbf{y}(\mathbf{x})\approx\ell\partial_{\mathbf{v}_{0}}\mathbf{y}(\mathbf{x}).$ (11)
Our first coarse-graining rule is to couple the cell-by-cell kinematics to the effective deformation via the constraints
$\partial_{\mathbf{u}_{0}}\mathbf{y}=\mathbf{R}\mathbf{u}(\theta),\qquad\partial_{\mathbf{v}_{0}}\mathbf{y}=\mathbf{R}\mathbf{v}(\theta),$ (12)
which match the righthand sides of (LABEL:eq:keyCoarseGrain10) and (LABEL:eq:keyCoarseGrain11). We can also eliminate the rotation field $\mathbf{R}$ and instead couple $\theta$ to $\mathbf{I}(\mathbf{y})$ by writing (12) equivalently as a metric constraint
$\mathbf{I}(\mathbf{y})=\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)$ (13)
for the the shape tensor in (8).
We now turn to our second coarse-graining rule. While most of the deformation of a soft mode is localized at the folds, the panels must bend slightly to accommodate a non-uniform actuation from cell to cell. This interplay constrains the effective theory by how neighboring unit cells with slightly bent panels fit together, leading to a fundamental relationship between the actuation $\theta$ and the second fundamental form of the effective deformation
@@
Fig. 4 illustrates the idea. Zoom into the local neighborhood of cells in a soft mode and identify the rotation of the base cell by $\mathbf{R}(\mathbf{x})$, as indicated, to be consistent with (LABEL:eq:keyCoarseGrain10) and Fig. 3. Given the slightly bent nature of all the panels in this description, the cells rotate relative to each other by a small amount. We quantify this kinematics at leading order by noticing that the rotation of the cell to the right and above the base cell satisfy
$\mathbf{R}(\mathbf{x}+\ell\tilde{\mathbf{u}}_{0})\approx\mathbf{R}(\mathbf{x})\big{(}\mathbf{I}+\ell(\bm{\omega}_{\mathbf{u}_{0}}(\mathbf{x})\times)\big{)},\qquad\mathbf{R}(\mathbf{x}+\ell\tilde{\mathbf{v}}_{0})\approx\mathbf{R}(\mathbf{x})\big{(}\mathbf{I}+\ell(\bm{\omega}_{\mathbf{v}_{0}}(\mathbf{x})\times)\big{)},$ (15)
@@
As a final bit of manipulation, this constraint can be rewritten using the implied identity $\partial_{\mathbf{u}_{0}}\big{[}\mathbf{R}\mathbf{v}(\theta)\big{]}=\partial_{\mathbf{v}_{0}}\big{[}\mathbf{R}\mathbf{u}(\theta)\big{]}$ from (12). Doing so eliminates the bending measures $\bm{\omega}_{\mathbf{u}_{0}}$, $\bm{\omega}_{\mathbf{v}_{0}}$ in favor of the second fundamental form $\mathbf{II}(\mathbf{y})$, giving an equivalent constraint in terms of the fields $\mathbf{y}$ and $\theta$:
@@ 2.3 Asymptotic analysis and an effective plate theory
The bulk of our effort in [64] goes into justifying the constraints in (13) and (17). We showed that every sufficiently smooth solution to these effective compatibility equations is the limit of a sequence of soft modes — origami deformations with negligible panel strain that approximate the effective deformation. The precise result is as follows: Let $\{\mathbf{x}_{i,j}^{\ell}\}\subset\mathbb{R}^{3}$ and $\Omega\subset\mathbb{R}^{2}$ be the origami pattern and its effective midsurface as above, so that the cell size $\ell$ is $\ll 1$ and the overall pattern has dimensions $|\Omega|\sim 1$. Then, for any sufficiently smooth $\theta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$ and $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$ that solve (13) and (17), there is a corresponding deformation of the origami vertices $\{\mathbf{x}_{i,j}^{\ell}\}\mapsto\{\mathbf{y}_{i,j}^{\ell}\}$ such that
$\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i+1,j}^{\ell}|-|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j}^{\ell}|}+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i,j+1}^{\ell}|-|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}$ (18)
$\qquad+\frac{|\mathbf{y}_{i,j}^{\ell}-\mathbf{y}_{i+1,j+1}^{\ell}|-|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j+1}^{\ell}|}{|\mathbf{x}_{i,j}^{\ell}-\mathbf{x}_{i+1,j+1}^{\ell}|}+\frac{|\mathbf{y}_{i+1,j}^{\ell}-\mathbf{y}_{i,j+1}^{\ell}|-|\mathbf{x}_{i+1,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}{|\mathbf{x}_{i+1,j}^{\ell}-\mathbf{x}_{i,j+1}^{\ell}|}=O(\ell^{2}),$
$|\mathbf{y}(\mathbf{x})-\mathbf{y}_{i,j}^{\ell}|=O(\ell)$
for all $(i,j)$, where $\mathbf{x}\in\Omega$ and $\mathbf{x}_{i,j}^{\ell}$ are corresponding points (see Fig. 3(a) and (c)). The first statement in (LABEL:eq:mainResult) says that the strain of every panel is at most $\sim\ell^{2}$. The second says that the origami deformation approximates the effective deformation.
In fact, we can say more. As another key result of [64], we derived an effective plate theory for parallelogram origami as the asymptotic limit of a large family of prototypical bar-and-hinge models of the origami [26, 33]. Section 2.2 of [64] explains the setup of our derivation in full detail. In brief, such models are composed of a sum of stretching, bending and folding elastic energies, denoted $E_{\text{tot}}^{\ell}(\{\mathbf{y}_{i,j}^{\ell}\})=E^{\ell}_{\text{str}}(\{\mathbf{y}_{i,j}^{\ell}\})+E^{\ell}_{\text{bend}}(\{\mathbf{y}_{i,j}^{\ell}\})+E^{\ell}_{\text{fold}}(\{\mathbf{y}_{i,j}^{\ell}\})$. The stretching elasticity is an energy that scales with the magnitude of the four panel strains in the first equation in (LABEL:eq:mainResult). The folding elasticity penalizes the deviations of each fold’s dihedral angle from one set by the origami’s reference configuration. Finally, the bending elasticity is modeled by adding “stiff” folds along the panel diagonals and measuring the deviation of these folds from flat under a deformation.
After scaling the moduli of the bar-and-hinge model by powers of $\ell$ such that “stretching stiffness” $\gg$ “bending stiffness” $\gg$ “folding stiffness”, we showed that the sequence of origami constructions in (LABEL:eq:mainResult) admits the following limit energy:
@@
where $\mathbb{K}(\theta)\in\mathbb{R}^{3\times 2\times 3\times 2}$ is an explicit fourth order tensor with major symmetry $[\mathbb{K}(\theta)]_{i\alpha j\beta}=[\mathbb{K}(\theta)]_{j\beta i\alpha}$ that depends kinematically only on the actuation field $\theta$ (see [64] for the full definition) [^4] In bar-and-hinge elasticity, there is always a question of which panel diagonal to choose for the stiff folds in the bending energy. We considered all possible choices in [64] and showed that the limiting plate theory is of the form in (19) for every case. However, the moduli $\mathbb{K}(\theta)$ depends on our choices for the panel diagonals. All else being equal, $\mathbb{K}(\theta)$ is stiffer if the longer panel diagonals are chosen and softer for the shorter ones. This result is consistent with the heuristic that shorter panel diagonals are the more physically sensible modeling choice. and $(\mathbf{II}(\mathbf{y}),\nabla\theta)^{T}\in\mathbb{R}^{3\times 2}$ denote the “bending” strains of the effective plate. [^5] As a heuristic for why $\nabla\theta$ is a bending strain, consider the bowtie example in Fig. 1(a). This origami deformation is effectively flat, which means the corresponding effective deformation in our theory satisfies $\mathbf{II}(\mathbf{y})=\mathbf{0}$. However, the panels are bending to achieve this configuration and thus storing some elastic energy. This energy is quantified by $\nabla\theta$ in our plate theory in (19). The set $A_{\text{eff}}$ denotes the class of admissible effective deformations and actuation fields describing a soft mode
@@
where $H^{k}$ is the space of square integrable functions whose $1^{\text{st}},\ldots,$ and $k^{\text{th}}$-order partial derivatives are also square integrable.